### v3

# Efficient Iterative Proximal Variational Inference Motion Planning

Abstract
Motion planning under uncertainty can be cast as a stochastic optimal control problem where the optimal posterior distribution has an explicit form. To approximate this posterior, this work frames an optimization problem in the space of Gaussian distributions by solving a Variational Inference (VI) in the path distribution space. For linear-Gaussian stochastic dynamics, we propose a proximal algorithm to solve for an optimal Gaussian proposal iteratively. The computational bottleneck is evaluating the gradients with respect to the proposal over a dense trajectory. We exploit the sparse motion planning factor graph and Gaussian Belief Propagation (GBP), allowing for parallel computing of these gradients on Graphics Processing Units (GPUs). We term the novel paradigm as the Parallel Gaussian Variational Inference Motion Planning (P-GVIMP). Building on the efficient algorithm for linear Gaussian systems, we then propose an iterative paradigm based on Statistical Linear Regression (SLR) techniques to solve motion planning for nonlinear stochastic systems, where the P-GVIMP serves as a sub-routine for the linearized time-varying system. We validate the proposed framework on various robotic systems, demonstrating significant speed acceleration achieved by leveraging parallel computation and successful planning solutions for nonlinear systems under uncertainty. An open-sourced implementation is presented at https://github.com/hzyu17/VIMP.
1 Introduction
Motion Planning is one core decision-making component in autonomous robotic systems [1, 2]. Given an environment, a start configuration and a goal configuration, a motion planner computes a trajectory connecting the two configurations. The trajectory optimization paradigm [3, 4] formulates the motion planning problem as an optimization over all admissible trajectories with dynamics and environment constraints. An ‘optimal’ plan can be obtained by solving the optimization program by minimizing certain optimality indexes, such as time consumption and control energy.
Uncertainties such as sensing [5], actuation noises, and external disturbances [6] arise and affect the quality of the motion plans in the execution phase. To reduce their impacts, motion planning under uncertainties takes into account the uncertainties in the planning phase in their formulations [7]. Stochastic optimal control and probabilistic robotics [8, 9] provide a principled framework to address this problem, where noise is explicitly modeled in robot dynamics, and a statistical optimality index is minimized over the trajectory distribution space.
Gaussian Process Motion Planning (GPMP) paradigm [10, 11] used the linear SDE dynamics model to formulate motion planning problems as probability inference, where the optimal solution is formulated as a posterior probability. The objective in GPMP is solving a trajectory that maximizes this posterior.
The solution obtained from GPMP is still deterministic. Variational Inference (VI) [12, 13] formulates the inference problem into an optimization by minimizing the Kullback–Leibler (KL) divergence between a proposal distribution and the target posterior. VI has been applied to solve for a trajectory distribution in the planning-as-inference formulation [14, 15, 16, 17]. Gaussian Variational Inference Motion Planning (GVIMP) [14, 15] showed satisfying performances in challenging tasks such as planning through narrow gaps [18]. The GVIMP formulation is related to entropy-maximized motion planning and reinforcement learning [19, 20, 21, 22].
Introducing distributional variables increases the dimension of the problem. Fortunately, in motion planning problems, the underlying probabilistic graph enjoys a sparsity pattern [23, 11] which is leveraged in algorithm designs to factorize the target posterior and reduce the time complexity [24, 15]. In this work, we study the proximal point GVI algorithm [25, 26] over the motion planning factor graph, where we show the updates in explicit form and that these gradient computations can be distributed onto the factor levels over the factor graph.
In terms of computation efficiency, similar to deterministic trajectory optimization paradigms [10, 11, 4], the collision-checking step of the algorithm on a dense trajectory is computationally demanding in GVIMP and represents the computational bottleneck [15]. In this work, we further leverage the sparse factor graph to distribute the collision-checking computations in parallel on a GPU. We also deployed the Gaussian Belief Propagation (GBP) [27, 28] algorithm for efficient marginal covariance computation. We term the algorithm the Parallel Gaussian Variational Inference Motion Planning (PGVIMP).
Most existing planning-as-inference paradigms [11, 29, 14] assume linear SDE as the underlying dynamics. Different from these works, we solve the problem for nonlinear stochastic differential equations (SDEs) [30], where we make the standard Gaussian transition assumption widely used in robotics applications [31, 32]. We then perform the Statistic Linear Regression (SLR) iteratively to obtain a linear time-varying (LTV) system in each iteration and perform PGVIMP on the obtained LTV system. This process integrates the sequential Bayesian principles into the GVIMP framework for motion planning tasks.
Related Works.
(a) Proximal Variational Inference. The proximal point algorithm and its stochastic version were first proposed in [25], where the KL-divergence was used as the proximal operator. The connection between the proximal-point and natural gradient descent algorithms was pointed out in [25, 13]. [26] generalized the proximal point variational inference algorithms by replacing the proximal term with general divergence functions. Recently, [33] proposed a proximal gradient algorithm by decomposing the KL-divergence objective into a potential and an entropy term. The above algorithms were applied directly to the joint posterior over all the variables without considering the factor graph structure of the posterior.
(b) Iterative approximations for nonlinear stochastic processes. The techniques based on iterative linearization to approximate nonlinear stochastic processes originated in the nonlinear state estimation literature. Related works are the Extended Kalman Filter (EKF), which leverages the Taylor approximation [34], statistical approximations based on moment matching [34], Laplace approximation [35], and cubature approximation [36]. In [37] the authors first introduced the use of Gauss-Hermite quadrature methods in the nonlinear estimation problems. To the best of the author’s knowledge, this work is the first to solve nonlinear planning problems by integrating iterative approximations and variational inference.
(c) Parallel Motion Planning. Parallelism and GPU acceleration have been applied to motion planners. The works [38, 39, 40, 41] proposed GPU-accelerated sampling-based motion planning algorithms. [42, 43] used GPU for fast collision-checking, which is a bottleneck in motion planning. The authors of [44] applied GPU acceleration in model predictive control algorithm for online control of high-DOF robotic systems. All these works dealt with collision-checking for deterministic states, while in our work, we solve collision-checking in a probabilistic sense, i.e., we compute an expected collision cost with respect to a proposal distribution over a dense trajectory.
In this work, we leverage this sparse factor and the parallel computing on GPUs to accelerate the computation of two key components in the GVI motion planning paradigm: the parallel computation of the collision likelihood and the computation of marginal covariances over the sparse factor graph.
Key Contributions. The contributions of this work are: (1) We proposed a KL-proximal variational inference algorithm for robot motion planning, where we show that the proximal gradient computations in the paradigm can be parallelized by leveraging a splitting structure and the underlying sparse factor graph; (2) A parallel computation implementation on GPU of the KL-proximal algorithms is proposed, where we leveraged parallel computation of the collision costs and Gaussian Belief Propagation (GBP) to compute marginal covariances efficiently; (3) We propose a sequential statistical linearization scheme for motion planning of nonlinear dynamical systems, where we use the parallel KL-proximal algorithm for linear systems as a submodule at each outer iteration.
Planning results are demonstrated on various robot models and comparison studies have been conducted to demonstrate the efficiency improvement.
This work builds upon the author’s prior work [14, 15], where we addressed the issue of computation efficiency, explored the KL-proximal GVI algorithm in the path distribution space for motion planning, and the applicability of the GVIMP approach to nonlinear dynamics.
This paper is organized as follows. Section 2 introduces the background knowledge. Section 3 introduces the KL-proximal GVI algorithm and its structure for motion planning problems. Section 4 introduces the sparse motion planning factor graph and marginal gradient updates for the KL-proximal algorithm. Section 5 introduces the iterative PGVIMP algorithm for nonlinear systems. The proposed framework is illustrated in Section 7 through numerical experiments, followed by a conclusion in Section 8.
2 Variational Inference Motion Planning (VIMP)
This section introduces some important preliminary knowledge of this paper, including motion planning under uncertainty and variational inference motion planning.
2.1 Planning under Uncertainties as Stochastic Control
We solve the motion planning under uncertainties using the stochastic control framework. For the nonlinear process
$dX_{t}=f(X_{t},u_{t})dt+g_{t}(X_{t})dW_{t},$ (1)
we minimize the following objective
$\!\!\!\!\min_{X(\cdot),u(\cdot)}$ $\mathcal{J}\triangleq\mathbb{E}\left\{\int_{0}^{T}\frac{1}{2}\|u_{t}\|^{2}+V(X_{t})dt+\Psi_{T}(X_{T})\right\}$ (2)
[…]
$dX_{t}=f_{0}(X_{t})dt+g_{t}(X_{t})dW_{t}.$ (3)
Gaussian assumptions. The state distribution of system (1) follows the Fokker-Planck (FPK) equations, leading to non-Gaussian distributions for the states under the nonlinearity (1). It is hard to solve the FPK for general nonlinearities , and Gaussian assumptions are one popular approximation assumption in the robotics applications, especially in the filtering and estimation [32, 31].
2.2 Gaussian Variational Inference Motion Planning (GVIMP)
Problem (2) admits an equivalent path distribution control formulation that will be introduced in this section.
(a) Control-inference duality. Denote the measure induced by the controlled linear process (7) as ${\mathbb{Q}}$, and the measure induced by the prior process (3) as ${\mathbb{P}}$. By Girsanov’s theorem [45], the following objective is equivalent to (2) [15, 46]
$\!\!\!\mathcal{J}$ $=\mathbb{E}_{{\mathbb{Q}}}[\log\frac{d{\mathbb{Q}}}{d{\mathbb{P}}}+\mathcal{V}]\propto\mathbb{E}_{{\mathbb{Q}}}[\log d{\mathbb{Q}}-\log\frac{e^{-\mathcal{V}}d{\mathbb{P}}}{\mathbb{E}_{{\mathbb{P}}}\left[e^{-\mathcal{V}}\right]}]$
$={\rm KL}\left({\mathbb{Q}}\parallel{\mathbb{Q}}^{\star}\right),$ (4)
[…]
$d{\mathbb{Q}}^{\star}\triangleq\frac{e^{-\mathcal{V}}}{\mathbb{E}_{{\mathbb{P}}}\left[e^{-\mathcal{V}}\right]}d{\mathbb{P}}\propto e^{-\mathcal{V}}d{\mathbb{P}}.$ (5)
Path space Variational Inference is formulated as [10, 15]
$q^{\star}=\underset{q_{\theta}\in\mathcal{Q}}{\arg\min}\;{\rm KL}\left(q_{\theta}\parallel{\mathbb{Q}}^{\star}\right)$ (6)
where $\mathcal{Q}\triangleq\{q_{\theta}:q_{\theta}\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})\}$ consists of the parameterized proposal Gaussian distributions.
(b) Linear-Gaussian System Case. When the stochastic process (1) reduces to linear time-varying stochastic system
$dX_{t}=(A_{t}X_{t}+a_{t}+B_{t}u_{t})dt+B_{t}dW_{t},$ (7)
the associated prior process is
$dX_{t}=(A_{t}X_{t}+a_{t})dt+B_{t}dW_{t}.$ (8)
Consider the above KL-minimization problem (4) with the time discretization
$\mathbf{t}\triangleq\left[t_{0},\dots,t_{N}\right],\;t_{0}=0,\;t_{N}=T.$ (9)
For linear Gaussian prior dynamics (8), the discrete-time path distribution $d{\mathbb{P}}$ over $\mathbf{t}$ is a Gaussian distribution $\mathbf{X}\sim\mathcal{N}({\boldsymbol{\mu}}|\mathbf{K})$. The discrete-time cost factor $e^{-\mathcal{V}}$ is defined as
[…]
then the un-normalized discrete-time optimal distribution to the problem (5) and is defined as
${\mathbb{Q}}^{\star}(\mathbf{X})\propto e^{-\frac{1}{2}\lVert\mathbf{X}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}-\sum_{i=0}^{N}V(X_{i})}\triangleq\tilde{q}^{\star}(\mathbf{X}).$ (10)
Under this discretization, the inference problem (6) becomes
$\begin{split}\underset{\mu_{\theta},\Sigma_{\theta}}{\min}\;\mathbb{E}_{q_{\theta}\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})}{\left[V(\mathbf{X})\right]}+{\rm KL}\left(q_{\theta}\parallel\mathcal{N}({\boldsymbol{\mu}},\mathbf{K})\right).\end{split}$ (11)
Define the negative log probability for the posterior as $\psi(X)\triangleq-\log\tilde{q}^{\star}(X)$. The Variational Inference objective is equivalent to
$\mathcal{J}(q_{\theta})={\rm KL}\left(q_{\theta}(\mathbf{X})\parallel\tilde{q}^{\star}(\mathbf{X})\right)=\mathbb{E}[\psi(X)]-\mathcal{H}(q_{\theta}),$
where $\mathcal{H}(q_{\theta})$ denotes the entropy of the distribution $q_{\theta}$, which promotes the robustness of the motion plan solution. We add an additional temperature parameter, $\hat{T}$, to trade off this robustness, leading to the objective of temperature [14]
$\mathcal{J}(q_{\theta})={\rm KL}\left(q_{\theta}(\mathbf{X})\parallel\tilde{q}^{\star}(\mathbf{X})\right)=\mathbb{E}[\psi(X)]-\hat{T}\mathcal{H}(q_{\theta})$ (12)
3 KL-Proximal GVI Algorithm For Linear-Gaussian Systems
In this section, we derive a proximal point Gaussian Variational Inference algorithm for the GVIMP problem for linear-Gaussian system formulation (11).
3.1 Proximal Point Gaussian Variational Inference
(a) KL-proximal point Variational Inference. The proximal point, or proximal minimization algorithm in the Euclidean space, is defined by solving the following optimization problem at each step
$x_{k+1}=\arg\min_{x}\mathcal{J}(x)+\frac{1}{2\beta_{k}}\lVert x-x_{k}\rVert^{2}_{2},$ (13)
where $\mathcal{J}$ is the original objective in a minimization problem, $\lVert x-x_{k}\rVert_{2}^{2}$ is a quadratic regularizer centered around the previous iterate $x_{k}$, and $\beta_{k}$ is the step size. The proximal point algorithm is shown to be more stable than vanilla gradient descent, as it can be interpreted as the backward Euler integration of the gradient flow [47, 48].
In the variational inference problems, the original objective is to maximize the ELBO, $\mathcal{L}$. The KL divergence is a better metric candidate than the Euclidean $2$-norm for the regularizer. Replacing the $2$-norm regularizer in (13) by the KL-divergence, we obtain at the KL-proximal point algorithm, which solves the following optimization at each step [49, 25]
$\delta\theta_{k+1}=\arg\max_{\theta}\mathcal{L}\left(x,\theta\right)-\frac{1}{\beta_{k}}{\rm KL}\left(q(x\mid\theta)\parallel q(x\mid\theta_{k})\right).$ (14)
(b) Splitting in Variational Motion Planning. To solve the proximal point iteration for general nonlinear ELBOs turned out to be challenging [26]. A commonly-encountered situation in variational inference is that the ELBO can be splitted into $2$ parts, $-\mathcal{L}=\psi_{e}+\psi_{d}$, where the gradients of $\psi_{e}\triangleq\mathbb{E}_{q_{\theta}}\left[\log\left(f_{e}(x)\right)\right]$ are easy to compute, sometimes are directly known in closed-form, e.g., for Gaussian priors and other conjugate models; and the computation of gradients of $\psi_{d}\triangleq\mathbb{E}_{q_{\theta}}\left[\log\left(f_{d}(x)\right)\right]$ is not direct. In order to obtain the updates efficiently, first-order approximations can be performed on $\psi_{d}$, resulting in proximal gradient algorithms [47]. This decomposition is also known as ‘splitting’.
Inspired by this, we derive the splitting structure for the variational inference in the path distribution space for motion planning problems for the first time. We decompose the motion planning ELBO into two different parts
$\begin{split}\!\!\!\mathcal{L}(x,\theta)&=\mathbb{E}_{q_{\theta}}\left[\log\tilde{{\mathbb{Q}}}^{*}\right]-\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}\right]\\ &=\!-\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}+\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}\right]\!-\!\mathbb{E}_{q_{\theta}}\left[V(\mathbf{x})\right],\end{split}$ (15)
where we use the two functions $\psi_{e},\psi_{d}$ to denote the two parts
$\psi_{e}(x,\theta)$ $\triangleq\mathbb{E}_{q_{\theta}}\left[\log q_{\theta}+\frac{1}{2}\lVert\mathbf{x}-{\boldsymbol{\mu}}\rVert_{\mathbf{K}^{-1}}^{2}\right],\;$ (16a)
$\psi_{d}(x,\theta)$ $\triangleq\mathbb{E}_{q_{\theta}}\left[V(\mathbf{x})\right].$ (16b)
The function $\psi_{e}$ is the negative likelihood of the prior cost-related term and the entropy, which is a conjugate model. The term $\psi_{d}$ is the cost related to collision avoidance, which is non-conjugate. Consider the Gaussian distribution $q_{e}\sim\mathcal{N}({\boldsymbol{\mu}},\mathbf{K})$, then $\psi_{e}$ can be written as the KL-divergence between $q_{\theta}$ and $q_{e}$ up to a constant as
$\psi_{e}(x,\theta)\propto\mathbb{E}_{q_{\theta}}\left(\log q_{\theta}-\log q_{e}\right)={\rm KL}\left(q_{\theta}\parallel q_{e}\right),$
which indicates that the gradients of $\psi_{e}$ with respect to $(\mu_{\theta},\Sigma_{\theta}^{-1})$ have closed-form. For the non-conjugate model, $\psi_{d}$, we approximate this term $\psi_{d}$ to the first order. The splitting and first-order approximation lead to the following KL proximal gradient update at each iteration.
(c) The KL-proximal gradients with splitting. After splitting, we derive the update step of the KL-proximal algorithm. Substituting the terms $\psi_{e}$ and $\psi_{d}$ in (16) into (14), we have
[…]
Taking the first-order gradients with respect to $(\mu_{\theta},\Sigma_{\theta}^{-1})$, we obtain the KL-proximal updates for them. The update rule is summarized in the following Theorem 1, whose proof can be found in A.
Theorem 1.
The KL-proximal updates for $(\mu_{\theta},\Sigma_{\theta}^{-1})$ are
$\!\!\!(\mathbf{K}^{-1}+\frac{\Sigma_{\theta_{k}}^{-1}}{\beta_{k}})\mu_{\theta_{k+1}}=-\nabla_{\mu_{\theta}}\psi_{d}+\mathbf{K}^{-1}{\boldsymbol{\mu}}+\frac{\Sigma_{\theta_{k}}^{-1}}{\beta_{k}}\mu_{\theta_{k}},$ (18a)
$\Sigma_{\theta_{k+1}}^{-1}=\frac{\beta_{k}}{\beta_{k}+1}(2\nabla_{\Sigma_{\theta}}\psi_{d}+\mathbf{K}^{-1}+\frac{\Sigma_{\theta_{k}}^{-1}}{\beta_{k}}).$ (18b)
Remark.
We remark that this update rule is accurate in the sense that we directly approximate the expected term $\psi_{d}(\theta)=\mathbb{E}_{q(\mathbf{x}\mid\theta)}\left[\|\mathbf{h}(\mathbf{x})\|_{\Sigma_{\mathrm{obs}}}^{2}\right]$ to the first-order, instead of approximating the integrand inside the expectation.
(d) Connection to Natural Gradient Descent (NGD). The NGD [24, 15] updates can be integrated into the proximal point algorithms with a symmetric KL-divergence being the proximal regularizer [13]. Specifically, the natural gradient descent update step with unit step size is
$\delta\theta=G(\theta)^{-1}\nabla_{\theta}\mathcal{L}(x,\theta),$
where $G(\theta)\triangleq\mathbb{E}_{q_{\theta}}\left[\nabla_{\theta}\log q_{\theta}\left(\nabla_{\theta}\log q_{\theta}\right)^{T}\right]$ is the Fisher information matrix. It is the solution to the following proximal point algorithm with the first-order approximation of the ELBO
$\delta\theta$ $=\arg\min_{d\theta}\;d\theta^{T}\nabla_{\theta}\mathcal{L}(x,\theta)+d\theta^{T}G(\theta)d\theta$ (19)
$\!\!\!\!\!\!\!\approx\arg\min_{d\theta}d\theta^{T}\nabla_{\theta}\mathcal{L}(x,\theta)+{\rm KL}^{\rm sym}(q(x\mid\theta)\parallel q(x\mid\theta+d\theta)),$
where the symmetric KL divergence ${\rm KL}^{\rm sym}$ between $2$ distributions $q_{1}$ and $q_{2}$ is defined as
${\rm KL}^{\rm sym}(q_{1}\parallel q_{2})\triangleq{\rm KL}(q_{1}\parallel q_{2})+{\rm KL}(q_{2}\parallel q_{1}).$
With a step size $\beta_{k}$, the NGD update is equivalent to changing the proximal operator in (17) from the KL divergence to the symmetric KL divergence.
4 Sparse Factor Graph For Marginal Computation
In the KL-proximal updates (18), computing the gradients $\nabla_{\mu_{\theta}}\psi_{d}$ and $\nabla_{\Sigma_{\theta}}\psi_{d}$ represents the computational bottleneck. This section introduces a paradigm that leverages the underlying sparse motion planning factor graph to distribute this computation onto the marginal levels.
4.1 Sparse Motion Planning Factor Graph
For clarity, denote $\psi(\mathbf{x})=V(\mathbf{x})$ in (16b). With some algebraic calculation, the gradients of $\psi_{d}$ are given by
$\nabla_{\mu_{\theta}}\psi_{d}(\theta)$ $=\Sigma_{\theta}^{-1}\mathbb{E}\left[\left(\mathbf{x}-\mu_{\theta}\right)\psi\right]$ (20a)
$\nabla_{\Sigma_{\theta}}\psi_{d}(\theta)$ $=-\frac{1}{2}\Sigma_{\theta}^{-1}\mathbb{E}[\psi]+\frac{1}{2}\Sigma_{\theta}^{-1}\mathbb{E}\left[(\mathbf{x}-\mu_{\theta})(\mathbf{x}-\mu_{\theta})^{T}\psi\right]\Sigma_{\theta}^{-1},$ (20b)
which can be broken apart into the calculation of the expectations of $\psi$
$\mathbb{E}\left[\psi\right],\;\mathbb{E}\left[\left(\mathbf{x}-\mu_{\theta}\right)\psi\right],\;\mathbb{E}\left[\left(\mathbf{x}-\mu_{\theta}\right)\left(\mathbf{x}-\mu_{\theta}\right)^{T}\psi\right].$
The motion planning factor graph is shown in the following Fig. 1. It contains two types of factors, the $f_{i,i+1}(x_{i},x_{i+1})$ represents the prior factor enforcing the kinodynamics constraints, and the $f_{i}(x_{i})$ are the collision factors that encourage the trajectory remain within obstacle-free regions.
Figure 1: The GVIMP factor graph.
This factorized structure can be leveraged to distribute the computation of $\psi_{d}$ on a GPU. Given the factor graph structure, the collision costs $\psi_{d}=\lVert h\left(x\right)\rVert^{2}_{\Sigma_{\rm obs}}$ can be factorized into
$\psi_{d}$ $=\psi^{d}_{0}(q_{0}),\dots,\psi^{d}_{N}(q_{N}),$
where the factorized potentials are defined as the collision cost on each individual state on the trajectory
$\psi^{d}_{i}$ $\triangleq\lVert h_{i}(q_{i})\rVert_{\sigma_{\rm obs}}^{i}.$
The factor-level marginal Gaussian variables are linearly mapped from the joint Gaussian
$q_{i}\sim\mathcal{N}\left(\mu_{\theta}^{i},\Sigma_{\theta}^{i}\right),\quad\mu_{\theta}^{i}=M_{i}\mu_{\theta},\quad\Sigma_{\theta}^{i}=M_{i}\Sigma_{\theta}M_{i}^{T}.$ (23)
Nest, we explore these sparsity patterns in the proposed Gaussian distributions.
4.2 Marginal updates
By Stein’s Lemma, we write (20a) and (20b) as
$\nabla_{\mu_{\theta}}\psi_{d}(\mathbf{x},\theta)$ $=\mathbb{E}\left[\nabla_{\mathbf{x}}\psi(\mathbf{x})\right],$ (24a)
$\nabla_{\Sigma_{\theta}}\psi_{d}(\mathbf{x},\theta)$ $=\frac{1}{2}\mathbb{E}\left[\nabla_{\mathbf{x}\mathbf{x}}^{2}\psi(\mathbf{x})\right].$ (24b)
Since we can decompose the $\psi(\mathbf{x})$ to marginal level
$\psi(\mathbf{x})=\sum_{n=1}^{N-1}\psi_{i}(\mathbf{x}_{i})=\sum_{i=1}^{N-1}\left\|\mathbf{h}\left(\mathbf{x}_{i}\right)\right\|_{\Sigma_{\mathrm{obs}_{i}}}^{2},$ (25)
where $\mathbf{x}_{i}=M_{i}\mathbf{x}$ is a subset of the variables relevant to the $i$th factor as defined in (23). Applying the factorization (25) to (24a), we obtain
$\mathbb{E}_{q}[\nabla_{\mathbf{x}}\psi(\mathbf{x})]=\sum_{i=1}^{N-1}M_{i}^{\mathrm{T}}E_{q_{i}}\left[\nabla_{\mathbf{x}_{i}}\psi_{i}\left(\mathbf{x}_{i}\right)\right].$ (26)
For the $i$-th factor, we can simplify the gradient and expectation to being over just $\mathbf{x}_{i}$, since $\psi_{i}$ only depends on $\mathbf{x}_{i}$. We have a similar result for (24b)
$\mathbb{E}_{q}\left[\nabla_{\mathbf{x}\mathbf{x}}^{2}\psi(\mathbf{x})\right]$ $=\sum_{i=1}^{N-1}M_{i}^{\mathrm{T}}\mathbb{E}_{q_{i}}\left[\nabla_{\mathbf{x}_{i}\mathbf{x}_{i}}^{2}\psi_{i}\left(\mathbf{x}_{i}\right)\right]M_{i}$ (27)
Apply Stein’s Lemma again to (26) and (27), we have
$\nabla_{\mu_{\theta}}\psi_{d}(\mathbf{x},\theta)$ $=\sum_{i=1}^{N-1}M_{i}^{\mathrm{T}}\nabla_{\mu_{\theta}^{i}}\psi^{d}_{i}(\mathbf{x}_{i},\theta_{i}),$ (28a)
$\nabla_{\Sigma_{\theta}}\psi_{d}(\mathbf{x},\theta)$ $=\sum_{i=1}^{N-1}M_{i}^{\mathrm{T}}\nabla_{\Sigma_{\theta}^{i}}\psi^{d}_{i}(\mathbf{x}_{i},\theta_{i})M_{i}.$ (28b)
The factorization of $\psi_{d}$ allows us to decompose the computation of its gradients across marginal distributions. Combined with parallel processing on a GPU, this structure significantly improves the computational efficiency of the algorithm.
4.3 Sparse Quadratures for Expectation Estimations
Our primary goal is to estimate the expectation (28) on the marginal levels. Gauss-Hermite quadrature approximations are widely used in filtering literatures [50, 32], as well as motion planning [14, 15]. Cubature approximations are accurate and efficient for lower dimensional problems. However, for higher dimensional problems, a straight-forward tensor product of the full-grid sigma points scale exponentially with the dimension, making it infeasible for high-dimensional problems [51].
Sparse-grid quadrature rules (Smoyak’s rules) [52, 53] ignore the cross terms among different dimensions to mitigate the curse of dimensionality, resulting in a quadrature rule with polynomial dimensional dependence. In this work, we adopt the sparse-grid Gauss-Hermite quadrature rules to estimate (28). To integrate a scalar function $\varphi(\mathbf{x})$
[…]
G-H quadrature methods consist of the following steps
  1. 1.
    Compute the $p$ roots of a $p\;th$ order Hermite polynomial, also denoted as sigma points: $\xi=[\xi_{1},\dots,\xi_{p}].$
  2. 2.
    Compute the weights: $W_{i}=\frac{p!}{p^{2}[H_{p-1}(\xi_{i})]},i=1,\dots,p.$
  3. 3.
    Approximate:
[…]
Lemma 1 (Smoyak sparse quadrature complexity).
For a $n-$variate function, the number of computations needed by a sparse grid quadrature that is exact for $p=2k_{q}-1$ total polynomial order is bounded by [53]
$\frac{e^{k_{q}}}{(k_{q}-1)!}n^{k_{q}}$
which has a polynomial dependence on the desired precision.
1
input : LTV system $\{(A_{t},a_{t},B_{t})\}_{t=1}^{N}$; Number of factors $L$; Upper bound of KL divergence $\epsilon$
output : Optimized trajectory distribution $\mathcal{N}(\mu_{\theta}^{*},\Sigma_{\theta}^{*})$.
2
3For $i=0,\dots,N$, compute the state transitional kernel $\Phi_{i,i+1}$, the Grammian $Q_{i,i+1}$, and the prior mean trajectory $\mu_{i}$.
4Compute prior precision $\mathbf{K}^{-1}$ and prior mean ${\boldsymbol{\mu}}$
5for $k=1,2,\ldots$ do
6
7      Compute marginals using GBP (23), (36). $\left\{q_{\ell}\sim\mathcal{N}\left(\mu_{\theta}^{\ell},\Sigma_{\theta}^{\ell}\right)\right\}_{\ell=1}^{L}\leftarrow\left(\mu_{\theta},\Sigma_{\theta}^{-1}\right)$
8      Parallel Computing the collision factors.
9      Mapping the gradients back to the joint level $\left(\nabla_{\mu_{\theta}}\psi_{d}(\mathbf{x}),\nabla_{\Sigma_{\theta}}\psi_{d}(\mathbf{x})\right)\leftarrow\left\{\left(\nabla_{\mu_{\theta}^{\ell}}\psi^{d}_{\ell}(\mathbf{x}_{\ell}),\nabla_{\Sigma_{\theta}^{\ell}}\psi^{d}_{\ell}(\mathbf{x}_{\ell}),M_{\ell}\right)\right\}_{\ell}$
10      Select the largest $\beta_{k}$ using bisection such that: $\mathrm{KL}(q_{\theta_{k+1}}\|q_{\theta_{k}})\leq\epsilon,\quad\Sigma_{\theta}^{-1}\succ 0$
11      Compute $\mu_{\theta_{k+1}}$ and $\Sigma_{\theta_{k+1}}^{-1}$ using (18) $\left(\mu_{\theta_{k+1}},\Sigma_{\theta_{k+1}}^{-1}\right)\leftarrow\left(\mu_{\theta_{k}},\Sigma_{\theta_{k}}^{-1},\nabla_{\mu_{\theta}}\psi_{d},\nabla_{\Sigma_{\theta}}\psi_{d},\mathbf{K}^{-1},{\boldsymbol{\mu}}\right)$
12
Algorithm 1 PGVIMP (LTV system).
1
input : Nonlinear system $\{(f_{t},g_{t})\}_{t=1}^{N}$, stepsize $\eta$; Nominal trajectory $\{(\bar{x}^{0}_{t},\bar{\Sigma}^{0}_{t})\}$.
output : Optimized trajectory distribution $\mathcal{N}(\mu_{\theta}^{*},\Sigma_{\theta}^{*})$.
2
3for $i=0,\dots,N$ do
4       Linearize the nonlinear system using SLR $\left\{\left(A^{i}_{t},a^{i}_{t},B^{i}_{t}\right)\right\}\leftarrow{\rm SLR}(f_{t},g_{t},\bar{x}^{i}_{t},\bar{\Sigma}_{t}^{i})$
5      Perform P-GVIMP (Algorithm 1) on the linearized system
6      $\left\{\left(\bar{z}^{i+1}_{t},\bar{\Sigma}^{i+1}_{t}\right)\right\}\leftarrow$ P-GVIMP$\left(A^{i}_{t},a^{i}_{t},B^{i}_{t}\right)$
7
Algorithm 2 i-PGVIMP (Nonlinear system).
5 Parallel GVIMP For Linear Time-Varying Systems
We first introduce a Parallel GVIMP algorithm on GPU for a linear time-varying (LTV) system.
5.1 Belief propagation for marginal covariances update
Computing the marginal updates (26) and (27) requires computing the marginal mean and covariance of the Gaussian distributions $q_{i}$. A major bottleneck is computing the marginal covariances. A brute force inverse computation of the joint covariance from $\Sigma^{-1}_{\theta}$ has the complexity $O((N\times n)^{3})$, which is a cubic function in $N\gg n$ in motion planning.
Leveraging the sparse factor graph, this section introduces the Gaussian Belief Propagation (GBP) [27, 28] to compute the marginals over the factor graph. Gaussian distribution $q\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})$ can be written in the canonical form
$q(x)\propto\exp(-\frac{1}{2}x^{\top}\Lambda_{\theta}x+\eta_{\theta}^{\top}x)$ (30)
where $\Lambda_{\theta}=\Sigma_{\theta}^{-1}$ denotes the precision matrix, and $\eta_{\theta}=\Sigma_{\theta}^{-1}\mu_{\theta}$ denotes the information vector
To simplify the computation, we can assume a shifted Gaussian distribution $r\sim\mathcal{N}(0,\Sigma_{\theta})$, since we only want to obtain the marginal covariance. Then $r(x)$ can be written as
$r(x)\propto\exp(-\frac{1}{2}x^{\top}\Lambda_{\theta}x)$ (31)
The precision matrix $\Lambda_{\theta}$ has such sparsity pattern
$\Lambda_{\theta}\!=\!\begin{bmatrix}\Lambda_{00}\!\!&\Lambda_{01}\!\!&&&&\\ \Lambda_{10}\!\!&\Lambda_{11}\!\!&\Lambda_{12}\!\!&&&\\ &\Lambda_{11}\!\!&\Lambda_{12}\!\!&&&\\ &&&\dots&&\\ &&&&\!\!\Lambda_{(N-1)(N-1)}\!\!&\Lambda_{(N-1)N}\\ &&&&\!\!\Lambda_{N(N-1)}\!\!&\Lambda_{NN}\end{bmatrix},$ (32)
and the Gaussian distribution (31) can be factorized into
$r(x)\propto\prod_{i=0}^{N}f_{i}(x_{i})\ \prod_{i=0}^{N-1}f_{i}(x_{i},x_{i+1})$ (33)
with
$f_{i}(x_{i})=$ $\exp\left(-\frac{1}{2}x_{i}^{T}\Lambda_{ii}x_{i}\right),$
$f_{i,i+1}(x_{i},x_{i+1})=$ $\exp\left(-\frac{1}{2}x_{i,i+1}^{T}\Lambda_{i,i+1}x_{i,i+1}\right),$
and
$x_{i,i+1}=\begin{bmatrix}x_{i}\\ x_{i+1}\end{bmatrix},\quad\Lambda_{i,i+1}=\begin{bmatrix}0&\Lambda_{i(i+1)}\\ \Lambda_{(i+1)i}&0\end{bmatrix}$
Gaussian Belief Propagation is an algorithm for calculating the marginals of a joint distribution via local message passing between nodes in a factor graph [28]. Since all of these factors are Gaussian, we can use this method directly.
Message passing on the factor graph can be divided into two types. The message passed from variables to factors, denoted as $m_{x_{i}\rightarrow f_{j}}$, and the message passed from factors to variables, denoted as $m_{f_{j}\rightarrow x_{i}}$. Messages can be obtained by
$m_{x_{i}\rightarrow f_{j}}$ $=\prod_{s\in N(i)\backslash j}m_{f_{s}\rightarrow x_{i}},$ (34a)
$m_{f_{j}\rightarrow x_{i}}$ $=\sum_{X_{j}\backslash x_{i}}f_{j}\left(X_{j}\right)\prod_{k\in N(j)\backslash i}m_{x_{k}\rightarrow f_{j}},$ (34b)
where $N(i)$ denotes all the neighboring factors of $x_{i}$, $f_{j}\left(X_{j}\right)$ denotes the potential of the factor $f_{j}$.
After computing all the messages in the factor graph, the beliefs of variables can be obtained by taking the product of incoming messages:
$b_{i}\left(x_{i}\right)=\prod_{s\in N(i)}m_{f_{s}\rightarrow x_{i}}$ (35)
since the messages here are also Gaussian distribution, we can obtain the belief parameters $\Lambda_{bi}$ by
$\Lambda_{b_{i}}=\sum_{s\in N(i)}\Lambda_{f_{s}\rightarrow x_{i}}$ (36)
The time complexity of inverting a $n\times n$ matrix is $O(n^{3})$, leading to a total complexity $O(N\times n^{3})$ for GBP.
5.2 Parallel Collision Checking on Dense Trajectory
One of the most computationally heavy modules for motion planning algorithms is the collision-checking module on a dense trajectory [4, 11]. This is also the case for variational motion planning algorithms [14, 15], where the collision-checking is required in a probabilistic sense, i.e., the goal is to compute the expected collision costs with respect to the proposal distribution, which is more challenging than computing a deterministic collision cost.
In our proposed method, we compute the expected collision costs by distributing the computation to each factor in the factor graph, as introduced in Section 4.2. Then, at marginal levels, we deploy the sparse GH quadratures to compute the marginal collision costs, as introduced in Section 4.3. After specifying the desired precision of this approximation, the quadrature method computes the sigma points and the weights, which are fixed throughout the experiment, and the expectation approximation is done by computing the matrix operations (29). Thus, one straightforward way of distributing the expectation approximation is by distributing the matrix multiplication (29), which is implemented in our proposed framework. The graphical demonstration of the parallel collision checking is shown in Fig. 2.
Figure 2: Parallel collision checking
5.3 PGVIMP Algorithm and Complexity Analysis
The PGVIMP algorithm is summarized in the following Algorithm 1. This algorithm serves as a sub-rouitne in motion planning for nonlinear systems by leveraging iterative linearization.
Complexity Analysis. We now analyze the time complexity of the proposed parallel GVIMP algorithm. The state dimension is $n$, and time discretization is $N+1$. The factor graph has $L=2N+3$ factors, the maximum dimension being $d=2n$. A serial computation of the nonlinear factors leads to a linear dependence on $N$, and the total complexity is $O(N\times n^{k_{q}})$, where $k_{q}$ is the desired precision in the GH-quadrature [15]. Experimentally, to obtain a precise estimate of the collision costs, $k_{q}\geq 6$ is required. The parallel nonlinear factor computation is of complexity $O(n)$, and the complexity of computing the marginal covariances for the tree-structured factor graph using GBP is of $O(N\times n^{3})$. The overall complexity of the algorithm is therefore reduced from $O(N\times n^{k_{q}})$ to $O(N\times n^{3})$, which is not dependent on the quadrature precision $k_{q}$.
6 Iterative P-GVIMP for Nonlinear Systems
Many robot dynamical systems are nonlinear. The P-GVIMP algorithm described in the previous sections constrained the system to be LTV. To handle the nonlinearity, this section introduces our proposed Iterative Parallel GVIMP Algorithm (i-PGVIMP).
6.1 Gaussian Assumptions for Nonlinear Transition
It is in general difficult to solve the FPK for arbitary nonlinear stochastic system (1). Gaussian assumption is popular in the robotics applications, especially in the filtering and estimation [32, 31]. It assumes the state distribution follows
$\frac{d\bar{X}_{t}}{dt}$ $=\mathbb{E}_{\mathcal{N}}\left[f(X_{t})\right],$ (37a)
$\frac{d{\rm Cov}(X_{t})}{dt}$ $=\mathbb{E}_{\mathcal{N}}[f(X_{t})(X_{t}-\bar{X}_{t})^{T}]+\mathbb{E}_{\mathcal{N}}[(X_{t}-\bar{X}_{t})f(X_{t})^{T}]$
$+\mathbb{E}_{\mathcal{N}}\left[g_{t}(X_{t})g_{t}(X_{t})^{T}\right],$ (37b)
where $\mathbb{E}_{\mathcal{N}}\left[\psi(\cdot)\right]\triangleq\int\psi(x)\mathcal{N}(x)$ denotes the expected values under Gaussian distributoins.
We adopt the Gaussian assumptions (37) and propose an iterative statistical linearization of the nonlinear system around a nominal trajectory. In this way, each iteration produces a LTV system from the linearization. We then perform the P-GVIMP algorithm 1 on the obtained linearized LTV system. The results are used in the next iteration for the SLR linearization. This process is performed until the linearization loop converges.
6.2 Iterative Statistical Linear Regression (SLR).
Statistical linear regression (SLR) [50] is a statistical method to approximate the nonlinear stochastic processes locally. The linearization results in an LTV system depending on the state at which the linearization is performed.
Assuming the closed-loop discrete nonlinear stochastic dynamics at time $t=t_{i}$ as $X_{i+1}=f_{\rm cl}(X_{i})$, then the goal of SLR is to find a linear model $(A_{i},a_{i})$
$\hat{X}_{i+1}=A_{i}X_{i}+a_{i}$
to minimize the statistical error, i.e.,
$\{A_{i},a_{i}\}=\arg\min\mathbb{E}\left[\lVert X_{i+1}-\hat{X}_{i+1}\rVert^{2}\right].$ (38)
Taking the first-order derivative with respect to $a_{i}$ and let it be zero, we arrive at
$a_{i}^{*}=\mathbb{E}\left[f_{\rm cl}\left(X_{i}\right)\right]-A_{i}^{*}\mathbb{E}\left[X_{i}\right].$ (39)
substituting (39) into (38), the gradients over $A_{i}$ is
$\left(f_{\rm cl}(X_{i})-\mathbb{E}\left[f_{\rm cl}(X_{i})\right]-A_{i}\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)\right)\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)^{T},$
letting which be zero leads to
$A_{i}^{*}=P_{yx}P_{xx}^{-1},$ (40)
where $P_{yx}$, $P_{xx}$ are defined respectively as
$P_{yx}$ $\triangleq\left(f_{\rm cl}(X_{i})-\mathbb{E}\left[f_{\rm cl}(X_{i})\right]\right)\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)^{T}$ (41)
$P_{xx}$ $\triangleq\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)^{T}.$ (42)
From (39) and (40), the computation of $(a_{i}^{*},A_{i}^{*})$ involves the computation of the expectations of a nonlinear function $\mathbb{E}\left[f_{\rm cl}(X_{i})\right]$, which in this paper we use the sparse GH-quadrature rules mentioned in Lemma 1. Specifically, at time $t_{i}$, for a given desired approximation precision, $k_{q}$, we compute the set of weights and sigma points $\{(W^{i}_{l},\xi^{i}_{l})\}_{l=1}^{n^{k_{q}}}$ for the sparse quadrature, and perform the quadrature approximation
$\mathbb{E}\left[X_{i+1}\right]=\mathbb{E}\left[f_{\rm cl}(X_{i})\right]\approx\sum_{l=1}^{n^{k_{q}}}W_{l}f(\sqrt{P_{xx}}\;\xi_{l}+\mathbb{E}\left[X_{i}\right]),$
which is then substituted into (39) and (40). This procedure is iteratively performed forward in time. The iterative Parallel GVIMP is summarized in Algorithm 2.
7 experiments
This section presents the numerical experiment results for the proposed method. The experiments are aimed to answer the following questions:
  • 1.
    Q1: How much efficiency do we gain by leveraging the parallel computations on GPUs, compared with serial computations on CPUs?
  • 2.
    Q2: The effectiveness of the KL-proximal algorithm in high-DOF robot motion planning tasks, and its comparison with the GPMP2 baseline;
  • 3.
    Q3: The effectiveness of the iterative PGVIMP algorithm in motion planning tasks for nonlinear dynamical systems.
To answer these questions, we first demonstrate the efficiency improvement brought by the parallel computing paradigm on a laptop GPU. Secondly, we show the planning results for two $7$-DOF linear robotics kinematic problems. Thirdly, we show the planning results for a linearized system and a full nonlinear planar quadrotor system. Finally, we present some ablation study results and comparisons with baseline methods.
7.1 Efficiency Improvement through Parallel Computation
The first set of experiments is conducted to showcase the improved computation efficiency by leveraging the parallel computations on a GPU. Efficiency improvement is gained from parallel collision cost computation and the computing of marginal covariances using GBP.
(a) Parallel collision-checking on a dense trajectory. To demonstrate the improvement in collision-checking efficiency on a dense trajectory, we first compare the time of computing the expected collision factors between the proposed GPU parallel computation and a serial implementation [15]. We made the comparison on $2$D / $3$D point robots and $7$-DOF WAM arm robots with linear time-invariant (LTI) system
$A_{t}=\begin{bmatrix}0&I\\ 0&0\end{bmatrix},~{}a_{t}=\begin{bmatrix}0\\ 0\end{bmatrix},~{}B_{t}=\begin{bmatrix}0\\ I\end{bmatrix}.$ (43)
$2$D Point Robot $3$D Point Robot $7$-DoF WAM
Serial $85.40$ ms $479.7$ ms $8010.5$ ms
Parallel $14.80$ ms $37.60$ ms $306.16$ ms
Improvement 82.67 % 92.16 % 96.18 %
Table 1: Implementation time comparison for Cost Expectation Estimations (quadrature precision degree: $k_{q}=10$ for 2D and 3D Point Robots, and $k_{q}=6$ for 7-DoF WAM Arm; time discretization $N=750$).
Figure 3: $7$-DOF WAM expected collision cost computation.
The results are shown in Tab. 1. On a dense trajectory (time discretization $N=750$), we achieved over $95\%$ improvement in the computation time compared with serial implementation on a CPU. Fig. 3 shows the computation time spent on the expected collision checking cost computation for different time discretizations. We observe that the parallel implementation is insensitive to the time discretization.
(b) GBP for computing the marginal covariances. Next, we demonstrate the efficiency improvement in computing the marginal covariances achieved through GBP. We compare with the brute-force way of computing the inverse of the precision matrix, $\Sigma_{\theta}^{-1}$, using a direct matrix inversion. The comparison results are in Tab. 2 and Fig. 4. We see a linear dependence on the number of time discretizations for GBP, and a polynomial ($N^{3}$) dependence of direct inversion. A $99.5\%$ improvement in implementation time is gained through GBP for a $7$-DOF WAM.
$2$D Point Robot $3$D Point Robot $7$-DoF WAM
Brute force $256.4$ ms $456.86$ ms $1234.04$ ms
GBP $2.36$ ms $3.01$ ms $6.04$ ms
Improvement 99.07% 99.34% 99.51%
Table 2: Implementation Time Comparison for Precision Matrix Inversion (Dimension of Precision: $5000$, $4998$, $4998$).
Figure 4: Time-Dimension curve for precision matrix inversion.
(c) The overall optimization time. Finally, we show the overall efficiency of the proposed paradigm by recording the total optimization time, and comparing with CPU serial implementation [14, 15]. The results are recorded in Tab. 3. We gained over $97\%$ improvement for the $7$-DOF planning problems.
$2$D Point Robot $3$D Point Robot $7$-DoF WAM
Serial $7.38$ s $27.07$ s $425.03$ s
Parallel $0.48$ s $1.12$ s $12.25$ s
Improvement 93.50 % 95.86 % 97.11 %
Table 3: Implementation Time Comparison for the whole optimization process (quadrature precision degree: $k_{q}=10$ for 2D and 3D Point Robots, and $k_{q}=6$ for 7-DoF WAM Arm; time discretization $N=750$).
Figure 5: Convergence of P-GVIMP inner-loop iterations for the linearized planar quadrotor system.
Figure 6: Convergence of the SLR iterations for the nonlinear planar quadrotor dynamical system.
Figure 7: Factorized and total costs in Fig. 5.
7.2 Planning Results for the 7-DOF Robot Arms
Next, we show the motion planning result for $2$ models of $7$-DOF robot arms, the WAM robot Arm [54], and the right arm of the PR2 robot [55]. We then visualize the results in the Robot Operating System (ROS) [56] with a high-fidelity robot model for both models integrated into the Moveit [57] motion planning package. Fig. 8 shows the planning results for the WAM robot arm, and Fig. 9 shows the results for the PR2 robot arm planning results. The experiment settings are listed in Tab. 7
7.3 Iterative PGVIMP for nonlinear dynamical system
We validate our proposed algorithm on a planar quadrotor system defined as
$\dot{X_{t}}=\begin{bmatrix}v_{x}\cos(\phi)-v_{z}\sin(\phi)\\ v_{x}\sin(\phi)+v_{z}\cos(\phi)\\ \dot{\phi}\\ v_{z}\dot{\phi}-g\sin(\phi)\\ -v_{x}\dot{\phi}-g\cos(\phi)\\ 0\end{bmatrix}+\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ 0&0\\ 1/m&1/m\\ l/J_{q}&-l/J_{q}\end{bmatrix}\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix},$ (44)
where $g$ is the gravity, $m$ represents the mass of the planar quadrotor, $l$ is the length of the body, and $J_{q}$ is the moment of inertia. $u_{1}$ and $u_{2}$ are the two thrust inputs to the system. In all our experiments, we set $m=1/\sqrt{2},l=\sqrt{2}$, and $J_{q}=1$. We linearize the system (44) around an initial nominal mean trajectory and perform the P-GVIMP algorithm on the obtained LTV system.
Iteration $1$ $3$ $10$ $20$
Norm difference $286.40$ $16.61$ $4.33$ $2.96$
Table 4: Norm of trajectory difference between two consecutive linearizations of the nonlinear dynamics.
Figure 8: Planning results for the WAM Robot for two tasks with $N=750$ states. The results are obtained from two algorithms: GPMP2 (blue) and KL-Proximal (Gray). For KL-Proximal, the settings are $T=2.0$ and $2.75$, $\Sigma_{obs}=20I$ and $19I$, The ratio between temperatures are $\hat{T}_{high}/\hat{T}_{low}=10$ and $15$, respectively, and $r+\epsilon_{sdf}=0.21$
Figure 9: Planning results for the PR2 Robot for two tasks with $N=750$ states. The results are obtained from two algorithms: GPMP2 (blue) and KL-Proximal (Gray). For KL-Proximal, the settings are $T=2.25$ and $2.0$, $\Sigma_{obs}=7.5I$ and $5I$, respectively, the ratio between temperatures is $\hat{T}_{high}/\hat{T}_{low}=10$, and $r+\epsilon_{sdf}=0.21$
(a) Convergence for LTV system. We first show the results for a linearized planar quadrotor system, using Algorithm 1. We show the convergence of the algorithm via a downsampled plot of the intermediate proposal Gaussian trajectory distributions in Fig. 5 with the low and high temperatures $\hat{T}$ in (12), and the corresponding cost evolutions, both for prior and collision factorized costs, and the total cost, in Fig. 7. After obtaining a collision-free trajectory with low temperatures, we switched to a high-temperature phase to put more weight on the entropy costs. The total costs keep decreasing throughout both processes.
WAMExp 1 WAMExp 2 PR2Exp 1 PR2Exp 2
P-GVIMP $10.26$s $9.48$s $14.03$s $11.86$s
GPMP2 $0.75$s $1.36$s $2.74$s $0.78$s
Table 5: The optimization time comparison for P-GVIMP and GPMP2 algorithms in the 7-DOF WAM and PR2 experiments with $n=750$ support states, averaged over 50 runs. For a 7-DOF robot, the GPMP2 algorithm has $14\!\times\!750=10,500$ optimization variables, while the KL-Proximal GVI has $14\!\times\!750+(3\!\times\!750\!-\!2)\times(14\!\times\!14)=443,608$ optimization variables.
(b) Planning results for the nonlinear planar quadrotor. Next, we consider the full nonlinear planar quadrotor system (44) using the iterative-PGVIMP planning, i.e., Algorithm 2. We compute the norm of the trajectory difference between consecutive iterations, and the iterations are stopped once the norm difference is below a threshold. The results are recorded in Tab. 4, and the planning results are shown in Fig. 6.
(c) More experiments with different settings. We did experiments for $4$ different settings. We have $1$ obstacle in the environment. The start and goal states in the $4$ experiments are specified in the following Tab. 8.
Fig. 11 shows the results of the nonlinear system of the planar quadrotor (44) for the $4$ experiment settings.
(d) Go through or go around a narrow gap? Robust motion planning through entropy regularization. As motion planning with obstacles is a naturally non-convex problem, the solution is multi-modal. In the decision-making for motion planning, the risk and robustness of the solution are taken into consideration in our formulation by introducing the entropy of $q_{\theta}$ into the objective. This section shows an experiment for the planar quadrotor to fly through a narrow gap [14]. Two motion plans are obtained for the same motion planning task, as shown in Fig. 10. One plan (go-through plan) is visually riskier than the other (go-around plan). Our formulation provides a computable metric to compare the total cost consisting of optimality and robustness, between the two plans. In Tab. 6, we compute the sum of the prior and the collision costs, the entropy costs, and the total costs. By introducing lower entropy costs, our method selects the plan with lower risk over a shorter but riskier plan.
Figure 10: Entropy regularized robust planning in a narrow gap environment.
Prior Collision MP Entropy Total
Go Through 138.96 1.26 140.22 1232.36 1372.58
Go Around 149.34 0.0369 149.38 1163.22 1312.6
Table 6: Comparing costs for two plans in Fig.10. The ‘MP’ is short for the sum of prior and collision costs, representing the optimality, and entropy represents the robustness.
(a) Experiment $1$.
(b) Experiment $2$.
(c) Experiment $3$.
(d) Experiment $4$.
Figure 11: Results of the i-PGVIMP for the nonlinear planar quadrotor dynamical system. $N=50$ support states are used, time span $T=4.0\sim 6.0$, radius $r+\epsilon_{sdf}=1.5$, $\Sigma_{obs}=6.0I\sim 7.5I$, low temperature $\hat{T}_{low}=1.0$, high temperature $\hat{T}_{high}=5.0$.
8 conclusion
In this work, we propose a distributed Gaussian Variational Inference approach to solve motion planning under uncertainties. The optimal trajectory distribution to a stochastic control problem serves as the target posterior in a variational inference paradigm. We leveraged this inference’s underlying sparse factor graph structure and proposed a distributed computation framework to solve the VI problem in parallel on GPU. Numerical experiments show the effectiveness of the proposed methods on an LTV system, and comparison studies demonstrated the computational efficiency.
Appendix A Proof of Theorem 1
Proof.
First take the gradient of the objectives in (17) with respect to $\mu_{\theta}$ and let it to be zero, we have
$0=\nabla_{\mu_{\theta}}\psi_{d}(\theta_{k})+\mathbf{K}^{-1}(\mu_{\theta}-{\boldsymbol{\mu}})+\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1}(\mu_{\theta}-\mu_{\theta_{k}}),$
from which we have the update for the mean as
$(\mathbf{K}^{-1}+\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1})\mu_{\theta_{k+1}}=-\nabla_{\mu_{\theta}}\psi_{d}(\theta_{k})+\mathbf{K}^{-1}{\boldsymbol{\mu}}+\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1}\mu_{\theta_{k}}.$
Since $\nabla_{\Sigma_{\theta}^{-1}}(\cdot)=-\Sigma_{\theta}\nabla_{\Sigma_{\theta}}(\cdot)\Sigma_{\theta}$, the necessary condition for the update rule for $\Sigma_{\theta}^{-1}$ can be derived by taking the gradients of the objectives in (17) with respect to $\Sigma_{\theta}$ and letting it to be zero, leading to
00 0 $=\nabla_{\Sigma_{\theta}}\psi_{d}(\theta_{k})+\frac{1}{2}(\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1}+\mathbf{K}^{-1}-(1+\frac{1}{\beta_{k}})\Sigma_{\theta_{k+1}}^{-1}),$
from which we obtain the update rule for the $\Sigma_{\theta}^{-1}$ as
$\Sigma_{\theta_{k+1}}^{-1}=\frac{\beta_{k}}{\beta_{k}+1}\left(2\nabla_{\Sigma_{\theta}}\psi_{d}(\theta_{k})+\mathbf{K}^{-1}+\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1}\right).$
Appendix B PGVIMP Algorithm Implementation Details
We present implementation details of the PGVIMP Algorithm 1.
(a) Step size selection. The proximal point algorithm is one type of gradient descent algorithm, which does not guarantee that the updated distribution at each step remains inside the distribution family $\mathcal{Q}$, especially with the step sizes being too large. For instance, the covariance updates in the Gaussian case does not guarantee that the updated covariances are always positive definite. We control the step size $\beta_{k}$ to enhance convergence efficiency by setting an upper bound on the KL divergence between the updated and current distributions. By solving the convex optimization problem
$\beta_{k}^{\star}\ =\ \max_{\beta_{k}>0}\ \beta_{k}$ (45)
[…]
we are able to find the largest step size $\beta_{k}$ that satisfies the constraints, ensuring both stability and efficient convergence. In practice, we use a bi-section algorithm to choose a suitable stepsize that keeps the next step distribution close to the current one.
Appendix C Experiment Settings
We record the settings for different experiments we conducted in this work in Tab. 7 and Tab. 8.
Start & Goal
WAMExp 1 $[-0.8,-1.70,1.64,1.29,1.1,-0.106,2.2]$
$[0,0.94,0,1.6,0,-0.919,1.55]$
WAMExp 2 $[-0.9,-1.7,1.34,1.19,0.8,-0.126,2.5]$
$[-0.7,-1.35,1.2,1.0,-0.7,-0.1,1.2]$
PR2Exp 1 $[-0.7,0.175,-0.785,-1.05,-2.88,-0.6,0.785]$
$[-0.174,0.785,-0.26,-0.785,1.57,-0.35,1.57]$
PR2Exp 2 $[-1.9,0.15,-2.18,-0.524,3.14,-0.524,-1.047]$
$[0,0.171,0,-0.96,3.14,-0.785,0]$
Table 7: $7$-DOF Robot Arm Planning Experiment Settings.
Start Goal
[…]
Table 8: Planar Quadrotor Planning Experiment Settings.